# Quantum Filtering at Finite Temperature

### v1

@@ III Quantum Filtering at Finite Temperature
We also introduce the process $Z_{t}^{\prime}=B_{t}^{\prime}+B_{t}^{\prime\ast}$. The $Z$ and $Z^{\prime}$ processes commute with each other, however, we note the nontrivial Ito table
@@
$dV_{t}=\bigg(\big[(n+1)L+nL^{\ast}\big]\otimes dZ_{t}-\sqrt{(n+1)n}\,(L+L^{\ast})\otimes dZ^{\prime}_{t}+K\otimes dt\bigg)V_{t},$ (38)
@@
$dA_{1,t}^{\ast}\,\Psi$ $=$ $\sqrt{n+1}\,dZ_{t}\,\Psi-\sqrt{n}\,dZ^{\prime}_{t}\,\Psi$
$dA_{2,t}^{\ast}\,\Psi$ $=$ $-\sqrt{n}\,dZ_{t}\,\Psi+\sqrt{n+1}\,dZ^{\prime}_{t}\,\Psi.$ (40)
@@
$dU_{t}\,\phi\otimes\Psi\equiv\bigg(\big[(n+1)L+nL^{\ast}\big]\otimes dZ_{t}-\sqrt{(n+1)n}\,(L+L^{\ast})\otimes dZ^{\prime}_{t}+K\otimes dt\bigg)U_{t}\,\phi\otimes\Psi.$ (41)
@@
$d\sigma_{t}(X)=\sigma_{t}(\mathcal{L}X)\,dt+\bigg\{(n+1)\sigma_{t}(XL+L^{\ast}X)+n\,\sigma_{t}(XL^{\ast}+LX)\bigg\}dY_{t}.$ (42)
@@
$V_{t}^{\ast}(X\otimes I)V_{t}$ $=$ $X\otimes I$ (43)
$+$ $\int_{0}^{t}V_{s}^{\ast}\big[X\big((n+1)L+nL^{\ast})\otimes I\big]V_{s}\otimes dZ_{s}+\int_{0}^{t}V_{s}^{\ast}\big[\big((n+1)L^{\ast}+nL)X\otimes I\big]V_{s}\otimes dZ_{s}$
$+$ $\sqrt{(n+1)n}\int_{0}^{t}V_{s}^{\ast}\big(\big[X(L+L^{\ast})+(L+L^{\ast})X\big]\otimes I\big)V_{s}\otimes dZ^{\prime}_{s}$
@@
$+(2n+1)\bigg((n+1)L^{\ast}+nL\bigg)X\bigg((n+1)L+nL^{\ast}\bigg)+(2n+1)(L+L^{\ast})X(L+L^{\ast})$
$-2\sqrt{(n+1)n}^{2}\bigg\{\bigg((n+1)L^{\ast}+nL\bigg)X(L+L^{\ast})+(L+L^{\ast})X\bigg((n+1)L+nL^{\ast}\bigg)\bigg\}.$
@@
$\mathbb{E}[V_{t}^{\ast}(X\otimes I)V_{t}|\mathfrak{Z}_{t}]$ $=$ $X\otimes I+\int_{0}^{t}\mathbb{E}\big[V_{s}^{\ast}\big[X\big((n+1)L+nL^{\ast})\otimes I\big]V_{s}|\mathfrak{Z_{s}}\big]\otimes dZ_{s}$ (45)
$+\int_{0}^{t}\mathbb{E}\big[V_{s}^{\ast}\big[\big((n+1)L^{\ast}+nL)X\otimes I\big]V_{s}|\mathfrak{Z_{s}}\big]\otimes dZ_{s}$
@@
$d\,\mathbb{E}[V_{t}^{\ast}(X\otimes I)V_{t}|\mathfrak{Z}_{t}]$ $=$ $\mathbb{E}\big[V_{t}^{\ast}\big[X\big((n+1)L+nL^{\ast})\otimes I\big]V_{t}|\mathfrak{Z_{t}}\big]\otimes dZ_{t}$ (46)
$+\mathbb{E}\big[V_{t}^{\ast}\big[\big((n+1)L^{\ast}+nL)\otimes I\big]V_{t}|\mathfrak{Z_{t}}\big]\otimes dZ_{t}$
@@
$d\pi_{t}(X)$ $=$ $\pi_{t}(\mathcal{L}X)\,dt+\bigg\{(n+1)\big[\pi_{t}(XL+L^{\ast}X)-\pi_{t}(X)\pi_{t}(L+L^{\ast})\big]$ (47)
$+n\big[\pi_{t}(XL^{\ast}+LX)-\pi_{t}(X)\pi_{t}(L+L^{\ast})\big]\bigg\}\,dI_{t}$
@@
$dI_{t}=dY_{t}-(2n+1)^{2}\pi_{t}(L+L^{\ast})\,dt.$ (48)
Proof. We first observe that the normalization factor satisfies $d\sigma_{t}(I)=(2n+1)\sigma_{t}(L+L^{\ast})\,dY_{t}$. From the Ito formula and (32), we see that
$d\frac{1}{\sigma_{t}(I)}=-(2n+1)\frac{\sigma_{t}(L+L^{\ast})}{\sigma_{t}(I)^{2}}\,dY_{t}+(2n+1)^{3}\frac{\sigma_{t}(L+L^{\ast})^{2}}{\sigma_{t}(I)^{3}}\,dt,$ (49)
@@
In the zero temperature case, we present the filter as a vector state valued process. For instance, Let $|\chi_{t}\rangle$ satisfy the stochastic differential equation $|d\chi_{t}\rangle=-(\frac{1}{2}L^{\ast}L+iH)|\chi_{t}\rangle+|\chi_{t}\rangle\,dY_{t}$ with $|\chi_{0}\rangle=|\psi_{0}\rangle$ a fixed normalized vector. Then $\sigma_{t}(X)=\langle\chi_{t}|X|\chi_{t}\rangle$ satisfies the quantum Zakai equation (42) with $n=0$ and initial condition $\sigma_{0}(X)=\langle\psi_{0}|X|\psi_{0}\rangle$. Likewise, $\|\chi_{t}\|^{2}\equiv\sigma_{t}(I)$ and we may write the filter as $\pi_{t}(X)=\langle\psi_{t}|X|\psi_{t}\rangle$ where $\psi_{t}=\chi_{t}/\|\chi_{t}\|$. This does not naturally extend to the $n>0$ case, however. Instead, we must start with (41) which leads to
$|d\Xi_{t}\rangle=\bigg(\big[(n+1)L+nL^{\ast}\big]\otimes dY_{t}-\sqrt{(n+1)n}\,(L+L^{\ast})\otimes dY^{\prime}_{t}+K\otimes dt\bigg)|\Xi_{t}\rangle.$ (51)
@@
$d\langle\Xi_{t}|X|\Xi_{t}\rangle$ $=$ $\langle\Xi_{t}|\mathscr{L}X|\Xi_{t}\rangle dt+\langle\Xi_{t}|\big[(n+1)L+nL^{\ast}\big]^{\ast}X+X\big[(n+1)L+nL^{\ast}\big]|\Xi_{t}\rangle dY_{t}$ (52)
$-\sqrt{(n+1)n}\,\langle\Xi_{t}|(L+L^{\ast})X+X(L+L^{\ast})|\Xi_{t}\rangle dY_{t}^{\prime}.$

### v2

@@ III Quantum Filtering at Finite Temperature
WE now come to a crucial point. Let us note the following nonzero quantum Ito table between the unprimed and primed noises:
$dB\,dB^{\prime}=dB^{\ast}\,dB^{\prime\ast}=dB^{\prime}\,dB=dB^{\prime\ast}dB^{\ast}=\sqrt{(n+1)n}\,dt.$ (33)
We introduce the process $Z_{t}^{\prime}=\frac{1}{i}(B_{t}^{\prime}-B_{t}^{\prime\ast})$. The $Z$ and $Z^{\prime}$ processes commute with each other, however, we note the nontrivial Ito table
@@
In principle, we could have chosen any quadrature $Z^{\prime}_{t}=e^{i\lambda}B_{t}^{\prime}+e^{-i\lambda}B_{t}^{\prime\ast}$ so that $Z$ and $Z^{\prime}$ are commuting (an therefore essentially classical) stochastic processes, but onlt the choice $\lambda=\pi/2$ leads to them being uncorrelated and therefore statistically independent.
@@
$dV_{t}=\bigg(\frac{1}{2n+1}\big[(n+1)L-nL^{\ast}\big]\otimes dZ_{t}-\frac{\sqrt{(n+1)n}}{2n+1}(L+L^{\ast})\otimes dZ^{\prime}_{t}+K\otimes dt\bigg)V_{t},$ (39)
@@
$dZ_{t}\,\Psi$ $=$ $\sqrt{n+1}\,dA_{1,t}^{\ast}\Psi+\sqrt{n}\,dA_{2,t}^{\ast}\Psi,$
$dZ_{t}^{\prime}\,\Psi$ $=$ $\frac{1}{i}\sqrt{n}\,dA_{1,t}^{\ast}\Psi-\frac{1}{i}\sqrt{n+1}\,dA_{2,t}^{\ast}\Psi,$ (41)
which may be inverted to give
$dA_{1,t}^{\ast}\,\Psi$ $=$ $\frac{\sqrt{n+1}}{2n+1}\,dZ_{t}\,\Psi+i\frac{\sqrt{n}}{2n+1}\,dZ^{\prime}_{t}\,\Psi$
$dA_{2,t}^{\ast}\,\Psi$ $=$ $\frac{\sqrt{n}}{2n+1}\,dZ_{t}\,\Psi-i\frac{\sqrt{n+1}}{2n+1}\,dZ^{\prime}_{t}\,\Psi.$ (42)
@@
$dU_{t}\,\phi\otimes\Psi$ $\equiv$ $\bigg(\big(\frac{n+1}{2n+1}L-\frac{n}{2n+1}L^{\ast}\big)\otimes dZ_{t}+i\frac{\sqrt{n(n+1)}}{2n+1}\,(L+L^{\ast})\otimes dZ^{\prime}_{t}$ (43)
$\qquad\qquad\qquad+K\otimes dt\bigg)U_{t}\,\phi\otimes\Psi.$
@@
$d\sigma_{t}(X)=\sigma_{t}(\mathcal{L}X)\,dt+\frac{1}{2n+1}\bigg\{(n+1)\sigma_{t}(XL+L^{\ast}X)-n\,\sigma_{t}(XL^{\ast}+LX)\bigg\}dY_{t}.$ (44)
@@
$V_{t}^{\ast}(X\otimes I)V_{t}$ $=$ $X\otimes I+\frac{1}{2n+1}\int_{0}^{t}V_{s}^{\ast}\big[X\big((n+1)L+nL^{\ast})\otimes I\big]V_{s}\otimes dZ_{s}$ (45)
$+$ $\frac{1}{2n+1}\int_{0}^{t}V_{s}^{\ast}\big[\big((n+1)L^{\ast}+nL)X\otimes I\big]V_{s}\otimes dZ_{s}$
$+$ $\frac{\sqrt{(n+1)n}}{2n+1}\int_{0}^{t}V_{s}^{\ast}\big(\big[X(L+L^{\ast})+(L+L^{\ast})X\big]\otimes I\big)V_{s}\otimes dZ^{\prime}_{s}$
@@
$+$ $\frac{1}{2n+1}\bigg((n+1)L^{\ast}-nL\bigg)X\bigg((n+1)L-nL^{\ast}\bigg)+\frac{n(n+1)}{2n+1}(L+L^{\ast})X(L+L^{\ast}).$
@@
$\mathbb{E}[V_{t}^{\ast}(X\otimes I)V_{t}|\mathfrak{Z}_{t}]$ $=$ $X\otimes I+\frac{1}{2n+1}\int_{0}^{t}\mathbb{E}\big[V_{s}^{\ast}\big[X\big((n+1)L+nL^{\ast})\otimes I\big]V_{s}|\mathfrak{Z_{s}}\big]\otimes dZ_{s}$ (47)
$-\frac{1}{2n+1}\int_{0}^{t}\mathbb{E}\big[V_{s}^{\ast}\big[\big((n+1)L^{\ast}+nL)X\otimes I\big]V_{s}|\mathfrak{Z_{s}}\big]\otimes dZ_{s}$
@@
$d\,\mathbb{E}[V_{t}^{\ast}(X\otimes I)V_{t}|\mathfrak{Z}_{t}]$ $=$ $\frac{1}{2n+1}\mathbb{E}\big[V_{t}^{\ast}\big[X\big((n+1)L-nL^{\ast})\otimes I\big]V_{t}|\mathfrak{Z_{t}}\big]\otimes dZ_{t}$ (48)
$-\frac{1}{2n+1}\mathbb{E}\big[V_{t}^{\ast}\big[\big((n+1)L^{\ast}-nL)\otimes I\big]V_{t}|\mathfrak{Z_{t}}\big]\otimes dZ_{t}$
@@
$d\pi_{t}(X)$ $=$ $\pi_{t}(\mathcal{L}X)\,dt+\bigg\{\frac{n+1}{2n+1}\big[\pi_{t}(XL+L^{\ast}X)-\pi_{t}(X)\pi_{t}(L+L^{\ast})\big]$ (49)
$-\frac{n}{2n+1}\big[\pi_{t}(XL^{\ast}+LX)-\pi_{t}(X)\pi_{t}(L+L^{\ast})\big]\bigg\}\,dI_{t}$
@@
$dI_{t}=dY_{t}-\pi_{t}(L+L^{\ast})\,dt.$ (50)
Proof. We first observe that the normalization factor satisfies $d\sigma_{t}(I)=\frac{1}{2n+1}\sigma_{t}(L+L^{\ast})\,dY_{t}$. From the Ito formula and (32), we see that
$d\frac{1}{\sigma_{t}(I)}=-\frac{1}{2n+1}\frac{\sigma_{t}(L+L^{\ast})}{\sigma_{t}(I)^{2}}\,dY_{t}+\frac{1}{2n+1}\frac{\sigma_{t}(L+L^{\ast})^{2}}{\sigma_{t}(I)^{3}}\,dt,$ (51)
@@
In the zero temperature case, we may present the filter as a vector state valued process. For instance, Let $|\chi_{t}\rangle$ satisfy the stochastic differential equation $|d\chi_{t}\rangle=-(\frac{1}{2}L^{\ast}L+iH)|\chi_{t}\rangle+|\chi_{t}\rangle\,dY_{t}$ with $|\chi_{0}\rangle=|\psi_{0}\rangle$ a fixed normalized vector. Then $\sigma_{t}(X)=\langle\chi_{t}|X|\chi_{t}\rangle$ satisfies the quantum Zakai equation (44) with $n=0$ and initial condition $\sigma_{0}(X)=\langle\psi_{0}|X|\psi_{0}\rangle$. Likewise, $\|\chi_{t}\|^{2}\equiv\sigma_{t}(I)$ and we may write the filter as $\pi_{t}(X)=\langle\psi_{t}|X|\psi_{t}\rangle$ where $\psi_{t}=\chi_{t}/\|\chi_{t}\|$. This does not naturally extend to the $n>0$ case, however. Instead, we must start with (43) which leads to
$|d\Xi_{t}\rangle=\bigg(\frac{1}{2n+1}\big[(n+1)L-nL^{\ast}\big]\otimes dY_{t}-\frac{\sqrt{(n+1)n}}{2n+1}\,(L+L^{\ast})\otimes dY^{\prime}_{t}+K\otimes dt\bigg)|\Xi_{t}\rangle.$ (53)
@@
$d\langle\Xi_{t}|X|\Xi_{t}\rangle$ $=$ $\langle\Xi_{t}|\mathscr{L}X|\Xi_{t}\rangle dt+\frac{1}{2n+1}\langle\Xi_{t}|\big[(n+1)L-nL^{\ast}\big]^{\ast}X+X\big[(n+1)L-nL^{\ast}\big]|\Xi_{t}\rangle dY_{t}$ (54)
$-\frac{\sqrt{(n+1)n}}{2n+1}\,\langle\Xi_{t}|(L+L^{\ast})X+X(L+L^{\ast})|\Xi_{t}\rangle dY_{t}^{\prime}.$